1. Given the Suns lifetime is about 10 billion years, estimate the lifetime of stars with (a)0.2-solar mass, (-5 x 10yr) (b) 3.0 solar mass, (4 x 10 yr) (c)10.0-solar mass x 10 yr) 2. A star is lying 20 pc from the Sun. The stars position on the sky is observed to move 5 arc sec per year relative to background stars. What is the stars transverse (perpendicular to the line of sight) velocity? If the observed red-shift is 0.001 percent (-Δλ/A), calculate the magnitude of the space velocity in m/s. (Take 1 pc 3 ly - 3x 10 m) (3.4 x 10 m/s) 3. A supernovas energy is often compared to the total energy output of the Sun over its lifetime (101 years). Assuming a constant luminosity of 4x 10 W, calculate the total energy output of the Sun. Compare this with the total energy ofa supernova emitted within months (about 1013 Joules). (-10H 10 26
Answer #1

a) ife time of star- -solar mass solar luminosity 0.2 ×1010 hje time of star =- 0.0004 life time of star 5.0x10 years b) 3.0 10, 75 life time of star =--x years life time of star =-4×10 years c) li fe time of star---x10 10 10000 lifetime of star=-1×10 years 3) ts minosity = 4 × 1026 W tine-1010yr = 10⑩ × 365×24× 60x60 sec so energy-lumin osity xtame energy = 4x1026W×100×365×24x60×60sec energy 1.26x104 J energy of sun-10; comparing the sun output energy in itslifetimeis 10times supernova energy emitted in amonth